OpenStudy (anonymous):

really need help! simply: (3m^3-1/2y)(3m^3-1/2y)

OpenStudy (anonymous):

*simplify, my bad

OpenStudy (ash2326):

We have \[(3m^3-\frac{1}{2} y)(3m^3-\frac{1}{2} y)\] Do you know the expansion of (a-b)^2?

OpenStudy (anonymous):

nooo? What is that?

OpenStudy (ash2326):

It's a standard relation always remember it \[(a-b)^2=(a-b)(a-b)=a^2-ab-ba+b^2=>a^2-2ab+b^2\]

OpenStudy (anonymous):

oh gosh

OpenStudy (ash2326):

We'll use that here \[(3m^2-\frac{1}{2}y)(3m^2-\frac{1}{2}y)={(3m^2-\frac{1}{2}y)}^2\] Did you not understand it?

OpenStudy (ash2326):

@becca94 ???

OpenStudy (anonymous):

sorry I was helping someone else really quick. umm... I kind of understand it. I'm not good at AlgII

OpenStudy (anonymous):

I've never heard of the standard relation and that's a lot in a specific order to remember.

OpenStudy (ash2326):

Ok \[(3m^2-\frac{1}{2}y)^2\] here \[a=3m^2\] \[b=-\frac{1}{2} y\] so using the formula, we get \[(3m^2-\frac{1}{2}y)^2= 9m^4-3m^2y+\frac{1}{4}y^2\]

OpenStudy (anonymous):

okay...I'm understand a bit better now

OpenStudy (ash2326):

Just remember \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

I'll write it down!

OpenStudy (ash2326):

Great:D